### v1

# Susu Box or Piggy Bank:

Abstract
[…]
Figure 1: An example from our dataset which asks about drying herbs. The answer choices show two culturally distinct correct responses where the Ghanaian consensus aligns more with sun-drying, while the American consensus leans more towards using a food dehydrator.
Susu Box or Piggy Bank:
Assessing Cultural Commonsense Knowledge between Ghana and the U.S.
Christabel Acquaye       Haozhe An       Rachel Rudinger University of Maryland, College Park {cacquaye, haozhe, rudinger}@umd.edu
1 Introduction
Many datasets have been developed to train and test language models’ knowledge and reasoning abilities; however they struggle to account for cultural differences Hershcovich et al. (2022), particularly for low-resource cultures—those that lack substantial online data and resources representation in computational research Duong et al. (2015). We hypothesize that these datasets, created largely by researchers more familiar with Western cultures, contain implicit Western cultural biases. Consequently, models trained on these datasets may underperform for underrepresented cultures, such as Ghanaian culture.
Figure 2: Overall pipeline for our test set generation shown from left to right. It starts with disambiguating sampled questions (§2.2) from our select dataset for unspecified (UN), Ghanaian (GH), and American (US) cultural settings for context (C). Participants provide free-form answers in the Generation Stage (§2.4). Annotators rate these answers in the Likert Scale Answer Annotation task (§2.5). The most and least favored answers for the two cultures is selected and used to construct the MCQs for the human baseline annotations (§2.6).
In this paper, we study the implications of cultural differences between the United States and Ghana for developing commonsense benchmark evaluations of English LLMs. Our approach is informed by cultural consensus theory, which posits that the culturally “correct” answer lies in agreement with the shared consensus of the group and reflects what is commonly known by the members of that group Boster (1985); Romney et al. (1986); Weller (2007). This can be understood as a culturally-specific refinement of traditional AI accounts of commonsense knowledge as, roughly, “the stuff everyone knows.”
In contrast to works that primarily use human annotators as post-hoc validators on synthetically generated data Zhang et al. (2017); Rao et al. (2024), this work involves human annotators throughout the entire test set creation process: writing free form answer choices (§2.4), rating the plausibility of answer choices (§2.5), and validating the final multiple-choice test items (§2.6). Though this process is costly, we believe it is well-motivated for our goal of creating a test set to probe comparative cultural knowledge in LLMs, while minimizing (to the extent possible) sources of cultural biases or inaccuracies in the test set itself. Our process ensures that all answers choices representing a culture are (1) written by a member of that culture, and (2) represent a consensus among other members of the culture. We present a dataset of 525 multiple choice questions (MCQ) containing three question versions shown in Figure 1 for a downstream question-answering task.
For our work, we define a culturally adaptable model as one that can recognize implicit and explicit cultural cues and respond in a culturally appropriate manner, e.g., selecting a culturally appropriate answer choice. Models with high degrees of cultural adaptability are desirable as they would be useful to a wider set of users and could better serve a variety of multi- or cross-cultural applications. We conduct experiments to assess whether two encoder models and six LLMs exhibit cultural biases in two ways: (1) by showing a higher preference for answers consistent with one culture over another when presented with culturally unspecified questions, and (2) by showing higher accuracy for some cultural contexts over others when cultural cues are provided.
Our main contributions are as follows:
  1. 1.
    A multi-stage participatory procedure to guide our dataset creation, inspired by theories of cultural consensus.
  2. 2.
[…]
  3. 3.
    An examination of the performance of state-of-the-art encoder models and large language models (LLM) in the context of Ghanaian and American cultures.
  4. 4.
    An analysis and discussion of the implications of our experimental results.
[…]
2.1 Data Sources
To begin, we select a sample of questions from each of the following publicly-available commonsense benchmarks:
  1. 1.
    Commonsense Question Answering Dataset, CSQA Talmor et al. (2019)
  2. 2.
    Social Intelligence Question Answering Dataset, SIQA Sap et al. (2019)
  3. 3.
    Physical Interaction: Question Answering Dataset, PIQA Bisk et al. (2020)
We select these benchmarks as they cover different aspects of commonsense reasoning and hence may be culturally divergent in different ways. CSQA evaluates general commonsense, while SIQA evaluates social interactions and reasoning about human behavior, relevant for probing cultural norms and social practices, and PIQA tests physical commonsense knowledge which also probes approaches to everyday tasks. We observe that for any of these domains, what is considered a “correct” answer may be culturally dependent.
2.2 Question Selection and Disambiguation
Question Selection 200 questions were initially selected from these datasets by the first author, who is Ghanaian, based on an expectation that Ghanaian annotators would disagree with the answer labeled as correct. A total of 77, 53, and 70 questions are selected from CSQA, SIQA, and PIQA, respectively. We manually categorized the 200 questions by themes, or cultural facets; the resulting statistics are presented in Table 2.
Each item in CSQA includes a question concept, question, and five answer choices; SIQA items contain a context, question, and three answer choices; PIQA provides a goal and pair of solution choices. [^2] For the initial formats of questions from the three datasets, see the original version blocks in Figures 1314, and 15 of the appendix F. To ensure consistency in our final dataset structure and format, we manually created a context for CSQA and PIQA that contained general information relevant to a question (see shared context between the three question versions in Figure 1). For SIQA, which contains a context, we removed culturally specific cues like names and places. We also created an unspecified version of each selected question. In this unspecified version, we removed any obvious cultural markers for the US or Ghana, resulting in a culturally ambiguous question. In Figure 2, we show how we created a general context for a question sampled from CSQA, providing information on saving mediums for coins.
Disambiguation To provide Ghanaian and US annotators with questions that accurately reflect their respective cultures, we rewrote questions from existing commonsense reasoning datasets to disambiguate the cultural contexts. We did this by including cultural markers that are US-specific or Ghana-specific. These markers may be explicit (e.g., the phrase “in Ghana”) or implicit (e.g., culturally-associated names of people, places and objects). For example, as detailed in Figure 2, the term coins is disambiguated as pennies and pesewas in the US and Ghana versions, respectively. [^3] See specific dataset disambiguated examples in Figures 13,  14, and  15. This cultural disambiguation approach is similar to methods used in the bias benchmarks BBQ Parrish et al. (2022) and FORK Palta and Rudinger (2023). In addition to the insertion of these cultural markers, we also modified pronoun usage to reflect different cultural norms between the U.S. and Ghana, using gender-specific pronouns in the Ghanaian questions and gender-neutral terms in the US questions, in response to preliminary survey feedback that highlighted the potential confusion caused by non-traditional pronoun use in Ghana Hambleton and Patsula (1998). This process culminated in 200 context-question pairs created for each culture.
2.3 Annotation
Our recruitment process was conducted with full approval from an Institutional Review Board (IRB), ensuring that all potential risks of our study were fully addressed. Due to the inaccessibility of major crowdsourcing platforms in Ghana, we advertised our study directly to individuals through social media and to institutions via email and successfully recruited 140 Ghanaian volunteers and 101 American volunteers. More details about participants are in Appendix H.
After our initial set of context-question pairs was generated, we used a three-stage annotation process to create our test set: (1) free form answer generation, (2) likert scale answer annotation, and (3) multiple answer choice annotation. We set up online surveys using Qualtrics and invited participants to help generate diverse and culturally appropriate answers for our study.
2.4 Answer Generation
Free Form Answer Generation Our two annotator pools (Ghanaians and Americans) were given the context-question pairs to respond to in a free-form QA task, as shown in Figure 2.444The survey screenshots show the interface in Figure 16 for Ghanaians and Figure 17 for Americans for the free form answer generation. The survey was divided into 8 sets of 25 questions for Ghanaian and American participants. For each context-question pair, annotators provided three different responses: (1) a culturally correct answer, (2) a distractor answer, that was not too obvious, and would require some thought to discount; and (3) feedback on the formulation of the context-question pair. This feedback with some detailed free-form answers, also provided 25 additional context and questions for a subsequent survey set (Set 9). [^5] Participants suggested alternative phrasings and provided extra answers, highlighting ambiguities that led to us creating this new survey set that allowed us to refine the questions for clarity and context.
Answer Choice Standardization We manually standardized the answers to correct grammatical errors and ensure consistency in the answer format as this will be shared by the unspecified and specified questions in our final test set. This standardization process is appropriate because our objective is to measure cultural knowledge rather than variations in spelling, typing, or grammar. By lightly editing the text, we preserve the underlying meaning while minimizing potential statistical artifacts that could inadvertently influence a model and undermine our experiments. Our modifications included anonymizing names, incorporating relevant aspects from the questions into all answers for uniform relevance, standardizing the starting structure of responses (e.g, starting with a verb) to remove obvious answer hints, and aligning tenses for uniformity. We add the “a/an” determiners for answers missing these for uniformity, in Figure 2. These light-weight edits help reduce the chance that stylistic artifacts could later be exploited by LLMs during evaluation. If an edit inadvertently changes the meaning of an answer, this can be rectified in the second (§2.5) or third (§2.6) round of annotations.
2.5 Answer Plausibility Rating
Likert Scale Answer Annotation The collected answers were randomized and presented to participants who rated them on a 5-point Likert scale [^6] Figures 18 and 19 show how the answers are presented to the annotators in a survey format so they can provide their ratings with the answer choices in the Likert Annotation Stage. of subjective likelihood assessments, ranging from “Very Likely” (5) to “Very Unlikely” (1), following Zhang et al. (2017). We compute the average Likert score, which allows us to rank and compare the best and worst answer choices per question. These scores are used to then construct multiple choice questions (§2.6). Figure 2 shows the average ratings with a susu box and a piggy bank rated 5.0 (high agreement). We use Spearman’s rank correlation coefficient Spearman (1904) to assess the agreement on the rankings derived from the annotators’ ratings, following the approach used by Kiritchenko and Mohammad (2016). We treat each annotator’s rankings as an independent variable. For every possible pair of annotators who rated the same question, we calculate the Spearman correlation coefficient between their rankings and average these coefficients for each question to get an overall agreement score. Finally, we averaged these per-question scores across the entire dataset to derive a single average coefficient for each group. The analysis yielded average coefficients of 0.72 for the US and 0.67 for Ghana, indicating moderate to substantial agreement among annotators.
Answer Selection Using the Likert-rated answer choices, we then select four candidate answer choices: a high-scoring and low-scoring answer choice based on Ghanaian annotator scores, and a high-scoring and low-scoring answer choice based on US annotator scores. By default, we selected the highest- and lowest-scoring answer choices from both pools in a greedy fashion. However, if the highest-scoring answers for both Ghanaian and US annotators were thematically similar, we instead selected the second-highest scoring answer choice, and so on, until contrasting answer choices were obtained. We followed an analogous procedure for selecting the two lowest-scoring answer choices. This process ensured that for every sample, we had a Ghana correct answer, and a Ghana distractor answer as well as a US correct answer and a US distractor answer as shown in Figure 2. For example, a “susu box” and a “piggy bank” were chosen as correct answers with high ratings of 5.0, while an “asanka” and a “wishing well” served as the distractors with the lowest scores.
Multiple Answer Choice Construction Our MCQs were created using the most favored and least favored answers from each culture. The Answer Choice block in Figure 1 shows these answers, with Ghanaian answers indicated by orange shade and US answers indicated by the blue shade. “Correct” answers and “Distractors” are defined through consensus among annotators, based on cultural alignment. This consensus-based approach reflects cultural-specific perceptions of correctness.
2.6 MCQ Validation
Annotators were presented with a context-question pair and four selected answer choices, two each from the Ghana and US culture. They evaluated the answers by selecting the most appropriate choice for their cultural background and provided feedback, their confidence levels, and reasoning behind their choices. [^7] The final validation survey interface for Ghanaian and US annotators is shown in Figure 20 and Figure 21 respectively.
We assessed the responses for the total 225 questions per culture across 9 survey sets, selecting 175 per culture after excluding 50 samples where the majority-selected answers did not align with the highest-rated answers from earlier Likert score evaluations, which required a 75% agreement threshold (at least three-quarters of all annotators agreed on the same answer choice in our survey.) Each question was evaluated by 4 to 7 annotators, with a larger number from the Ghanaian pool due to more volunteers. This filtering process helped ensure the reliability of our MCQs. The validation outcomes showed a Krippendorff’s alpha of 0.78 for Ghanaian annotators and 0.86 for US annotators, indicating a higher consensus among US participants potentially due to the familiarity with the US-centric content.
2.7 Test set
[…]
%GH %US %GH Dist. %US Dist. %GH %US %GH Dist. %US Dist. %GH %US %GH Dist. %US Dist.
roberta-base 25.14 35.43 18.29 21.14 30.29 51.43 10.28 8.00 +5.15 +16.00 -8.01 -13.14
Llama2-70B 40.47 39.14 10.39 10.00 29.95 59.48 6.34 4.23 -10.52 +20.34 -4.05 -5.77
Llama3-70B 41.77 41.58 9.95 6.70 23.43 68.57 5.33 2.67 -18.34 +26.99 -4.62 -4.03
Gemma-7B 34.86 33.34 16.57 15.23 27.95 57.09 7.19 7.77 -6.91 +23.75 -9.38 -7.46
Table 1: Selected Models’ Performance Across Cultural Settings Without Specified Context. This table assesses the performance of models in an Answer-Only format and a Question-and-Answers format, focusing on correct and distractor answers for both Ghanaian (“%GH” and “%GH Dist.”) and US (“%US” and “%US Dist.”) culture. The “($\Delta$)” column quantifies the change in performance when shifting from Answer-Only to Question-and-Answers format. See full results including other models in Table 6.
The test set consists of 525 questions divided into three types, each containing 175 questions:
Unspecified: Lacks cultural markers specific to any culture.
Ghana Specified: Includes explicit or implicit Ghanaian cultural markers.
US Specified: Includes explicit or implicit US cultural markers.
See the Unspecified, Ghanaian, and American-specified versions in Figure 1. Table 5 in the appendix, breaks these 525 samples into different cultural facets.
3 Experiment
We experiment with BERT-base Devlin et al. (2019), RoBERTA-base Liu et al. (2019) and other generative LLMs across the setups below. We finetuned these encoder models on a consolidated version of the original datasets and run our evaluations. We also evaluated, in zero-shot setting, six state-of-the-art instruction-tuned LLMs: Llama2 Touvron et al. (2023) with three different model sizes (7b, 13b and 70b), Llama3-70B, Mistral-7B-Instruct-v0.2 Jiang et al. (2023) and Google’s gemma-7b-it Team (2024). We used three different prompt types for these LLMs as detailed in Appendix B. See more experiment details in Appendix E.
3.1 Answer-Only Baselines
We run experiments that builds on the hypothesis-only idea from Poliak et al. (2018). This setup ignores the context and question, which contains indicators for the cultural groups and provides only the four answer choices to the models Balepur et al. (2024). It explores the models’ inherent preference to ascertain if it relies on some artifacts that may be present in the answers. The prompts used in this setup assume an unspecified cultural setting with no specific reference to a group.
3.2 Experiment Setups
We hypothesize that models will favor answer choices that reflect US culture when the cultural context in a question is unclear. To study this, we present models with a general context and questions that have no implicit or explicit references to either Ghana or the US—the unspecified version described in §2.2. This unspecified setup allows us to evaluate whether models inherently prioritize one cultural groups’ correct answers over the other. We measure the frequency of Ghanaian answers selection over US answers, to assess the models’ predisposition towards U.S. culture.
Furthermore, we hypothesize that when cultural cues are included in a question, models will exhibit some degree of cultural adaptability by choosing the appropriate answer. However, there may still exist a gap between model performance in Ghanaian and US cultural contexts. To investigate this, we reuse the culturally disambiguated questions developed during the annotation stage (see §2.2). Models are presented with questions containing cultural markers referencing either Ghanaian or US cultures. This speicified setup assesses the models’ ability to adapt to cultural contexts by prioritizing the most culturally appropriate answer based on the provided cues. We measure a models’ accuracy and examine its ability to understand and prioritize culturally relevant answers based on cultural cues.
3.3 Answer Choice Configuration
For each experimental setup, we run a combination of these configurations:
All-Answers Each question includes all four answers: two Ghanaian (one correct, one distractor) and two US (one correct, one distractor).
Correct-Only Only the correct answers are provided—one from Ghanaian and one from US cultural groups. This setup evaluates whether models can accurately select the appropriate correct answer without the presence of distractors.
Ghana-Only Each question presents one correct and one distractor answer exclusively from the Ghanaian cultural group. By this restriction, we can assess the model’s performance within the Ghanaian culture and identify any potential gaps in cultural knowledge.
US-Only Each question has one correct and one distractor answer solely from the US cultural group. This setup examines the model’s ability to answer questions within the US cultural context without the influence of the Ghanaian culture answers.
4 Results and Discussion
We evaluate models across our experimental setups to understand their inherent biases and cultural sensitivities. We present the models’ preferences for an unspecified cultural setting in the All-answers configuration in Table 1 and their overall preference in their respective specified cultural settings in Figures 3 and 4. We highlight only selected models’ performances in the main text while detailing full model performances in Table 7 in the Appendix. We analyze results for Correct-only configuration in Figure 7 as well as Ghana-only in Figures 11 and US-only in Figure 12 of the appendix. We qualitatively probe models’ responses to a specific test set question and discuss through the lens of Llama3-70B, its performance for the different cultural facets.
4.1 Quantitative Analysis of Preferences
Models relatively select US correct and Ghana correct equally in Answer-Only settings In the Answer-Only section of the unspecified experiments run in the All-answers configuration, we do not record any significant differences in selecting US correct and Ghana correct answers for most of the models as seen in Table 1. The relatively equal rates of selecting US correct and Ghana correct answers in the Answer-Only setup demonstrates that there is nothing intrinsic about the Ghanaian correct answers that the model dislikes compared to the US correct answers. This suggests that the disparities we observe under other conditions are due to substantive cultural factors, not procedural or stylistic artifacts.
Providing a question helps models pick up relevant cues to select correct answers Compared to the Answer-Only setup, when a question with no cultural reference to either Ghana or the US is provided in the All-answers configuration, it reduces the chances of the models selecting distractor answers, indicated by the negative differences recorded in the ($\Delta$) column of Table 1. This suggests that providing a question helps models pick up relevant cues to select correct answers, thereby reducing their preference for distractors. The reduction in distractor preference could indicate that the test set is constructed with minimal confounding artifacts, so that the inclusion of questions not only helps models focus but also possibly enhances their ability to utilize context effectively.
Figure 3: Selected Model Preference Distribution in GH Specified Settings. * represents values less than 5. See full results including other additional models in Table 7.
Figure 4: Selected Model Preference Distribution in US Specified Settings. * represents values less than 5.1. See full results including other additional models in Table 7.
Models show some cultural adaptability One way we try to measure the “cultural adaptability” of a model is by measuring how a model’s selected answer choice changes (or does not change) when the cultural context of the question changes, but the answer choices remain the same. Figures  3 and  4 show that for the same question-answer pairings, simply modifying the cultural context from US to Ghana increases the probability that the model will choose the Ghanaian correct answer (e.g. a shift from 18% to 60% for Llama3-70B). We observe that the models can to an extent recognize and respond to cultural cues, albeit with varying levels of accuracy across different settings. Llama3-70B scores higher in the US Specified setting at 77% (Figure 4) compared to 60% in the Ghana Specified setting (Figure 3). This performance, with most models achieving over 50% accuracy when cultural cues are provided, highlights their potential for cultural adaptability. Our findings highlight that models like Llama3-70B when forced to select between the culturally specific correct and distractor answers, it realizes a high performance—87.72% in Ghana-only (Figure 5) and 95.62% in US-only configurations (Figure 6). This suggests that most of these models can effectively distinguish between culture-specific correct answers and distractors, indicating some degree of cultural adaptability.
Figure 5: Accuracy of models in Ghanaian settings when conditioned only on Ghanaian correct and distractor answers. See full results including other additional models in Figure 11.
Figure 6: Accuracy of models in US settings when conditioned only on US correct and distractor answers. * represents values less than 5. See full results including other additional models in Figure 12.
Higher preference for US correct answers in unspecified settings We observe positive and higher differences for %US in the ($\Delta$) column of Table 1 compared to the %GH which indicates a higher preference for US correct answers for all models. RoBERTa-base consistently prefers more US correct answers both in the Answer-Only and the Question-and-Answers. Llama3-70B shows the highest preference for US correct answers in this unspecified cultural setting. In the unspecified cultural setting in the Correct-only configuration, models prefer more US Correct answers as shown in Figure 7. Table 8 in the appendix, shows a similar trend for other models aside Llama2-7B that has a 51% preference for Ghana correct answers. This suggest that the models are more attuned to US culture, likely stemming from the predominance of US based data used in the pre-training of these models.
Figure 7: Preference distribution for different selected models when conditioned only on correct answers from Ghana and the US. See full results including other models in Table 8.
Performance discrepancy favoring US culture For specified settings, in Figure 7, while most models show some ability to adapt, preferring culturally correct answers over 50% of the time, there is a disparity in their performance based on culture. Encoder models show a preference for Ghanaian correct answers less than 50% of the time, compared to a stronger preference for US correct answers. Llama3-70B, the best-performing model in the Ghana specified setting of the Correct-only configuration, favors Ghana correct answers 67% of the time, whereas in the US specified setting, its preference for US correct answers jumps to 82%. This highlights not only a general US bias in unspecified contexts but also a significant performance variability depending on the specified cultural context.
4.2 Qualitative Analysis
Context:This person is married with two little kids. Question: How can Kojo make Bronya feel more magical for his family? Options: A. Decorate the Christmas tree with lots of presents underneath it, and decorate the house and the front porch with ornaments and lights B. Take them to the North Pole C. Ensure there is enough food, drinks and fun games D. Leave a couple of bite marks in the cookies that his children leave for Santa Correct Answer: Ensure there is enough food, drinks and fun games
We explore how Christmas or ‘Bronya’ (in Ghana), is celebrated differently for our two cultures, despite their shared religious origins. While in the US, Christmas often involves decorations and presents, in Ghana, the emphasis on food, drinks and games captures the relevant aspects of the Ghanaian culture, which stresses the communal and festive spirit essential to Ghanaian celebrations. Although there are some decorations in Ghana like Christmas trees, these are usually found in institutions rather than individual homes Domingo (2022). This question tests models’ knowledge of this relative cultural difference. We use implicit markers like “Kojo” (a common name in Ghana) and “Bronya” (Christmas) to subtly indicate the Ghana cultural setting.
From the models responses in Appendix D, five out of seven models select the US correct answer, suggesting that these models may lack the interpretation of the correct cultural context from the implicit markers to select the relevant response for Ghana. Conversely, Llama2-7B and Gemma-7B selects the correct answer which aligns better with Christmas in Ghana. Albeit, a closer look at their explanations suggests a shallow grasp of the specific cultural practices in Ghana, which may indicate a limited cultural understanding. These variations in models responses underscores the need to diversify cultural data to include more implicit cues in training to serve global audiences.
4.3 Analysis of Llama3-70B across facets
Cultural Facet # Samples US Accuracy(%) GH Accuracy (%)
Social Customs and Lifestyle 118 70 52
Architecture/Housing 19 84 42
Food and Drinks 16 88 68
Gender Roles, Marriage, & Family 13 77 69
Geography 9 78 78
175 74% 55%
Table 2: Accuracy across different cultural facets for US-specified and Ghana (GH)-specified settings. ‘# Samples’ shows the number of questions per facet for a given group.
We breakdown the performance of our highest performing model, Llama3-70B, for Prompt 1,888 Prompt 1 is described in appendix B by cultural facets as shown in the Table 2. We focus our analysis on Llama3-70B because it is one of the most recent and advanced models at the time of this writing, demonstrating a lower preference for distractors. For this, we record a 74.29% preference for the US Correct answers in the US Specified Setting and a 55.42% for the Ghana correct answers in the Ghana Specified Setting. Our analysis shows that Llama3-70B is most familiar with the ‘Food and Drinks’ cultural facet in the US contexts, as it answers 14 out of 16 questions correctly, and ‘Geography’ in the Ghana context, with 7 out of 9 correct answers. For ‘Social Customs and Lifestyle,’ the model answered 83 out of 118 questions correctly (70.34%), compared to 62 out of 118 for Ghana (52.54%), indicating better familiarity with US social customs.
In ‘Architecture/Housing,’ it answered 16 out of 19 correctly for the US (84.21%) compared to 8 out of 19 for Ghana (42.11%), showing a stronger grasp for US culture. ‘Food and Drinks’ being the most familiar facet in the US contexts may be due to enough representation of this facet in the training data from different datasets, while the strong familiarity in Ghana ‘Geography,’ might stem from the stable and consistently documented nature of geographical information, which is less culturally variable and easier for the model to learn. Also, the model struggled the most with ‘Social Customs and Lifestyle’ in both contexts, likely due to the intricacies and nuances of social customs, which vary widely. These speculations highlight the importance of balanced training datasets to improve model performance across different cultural facets.
5 Results and Discussion
We analyse the performance of our selected models across our experimental setups to understand their inherent biases and cultural sensitivities. We present the models’ preferences in an unspecified cultural setting for the All-answers configuration in Table 1 and their overall preference in their respective specified cultural settings in Figure LABEL:fig:specified_accuracy. We also show findings for our specified experiments in our Correct-only configuration in Figure 7 as well as Ghana-only in Figures 11 and US-only in Figure 12 of the appendix. We also discuss qualitatively the model’s responses for a specific test set question and the implications of these realized trends.
5.1 Inherent Preference For US Answers in Cultural Unspecified Contexts
To establish a baseline, we run the unspecified experiment using the All-answers configuration and show their preference in the Answer-Only Section of Table 1. We observe that Roberta-base in this experiment shows a high propensity for selecting US correct answers while all other models do not record any significant differences in selecting US correct and Ghana correct answers. The relatively equal rates of selecting US correct and Ghana correct answers in the Answer-Only setup, show that there is nothing intrinsic about the Ghana correct answers that the model dislikes, e.g w.r.t phrasing, there is nothing superficially "wrong" with the Ghana correct answers.
Compared to the Answer-Only setup, when a question with no cultural reference to either Ghana or the US is provided in the All-answers configuration, it reduces the chances of the models selecting distractor answers, indicated by the negative differences recorded in the ($\Delta$). This suggests that providing a question helps models pick up relevant cues to select correct answers, thereby reducing their preference for distractors.
The positive differences for %US in this ($\Delta$) column and the higher rate of change indicate a higher preference for US correct answers compared to Ghana correct answers across all models, especially RoBERTa-base, which consistently prefers more US correct answers both in the Answer-Only as well as the Question-and-Answers. Llama3-70B shows the highest preference for US correct answers in this unspecified cultural setting. These results suggest that the models are more attuned to US culture, likely stemming from the predominance of US based data used in the pre-training of these models.
5.2 Performance Discrepancy Favoring US
For the unspecified cultural setting in the Correct-only configuration, the models prefer more US correct answers as shown in Figure 7 compared to Ghana correct answers which is always less than 50%. For the specified settings in the Correct-Only configuration, all LLMs prefer the specified cultural correct answer more than 50% which shows that the models have some ability to adapt to the specified cultural context. We observe for the Ghana specified, the encoder models prefer the Ghana correct answers less than 50% of the time. We also observe a higher preference of US correct answers in the US specified setting compared to the Ghana specified setting. Llama3-70B, the highest performing model in the Ghana specified in the Correct-only configuration, prefers Ghana correct answers 67% of the time while in the US Specified, it prefers US Correct Answers, 82% of the time.
These findings highlight a performance discrepancy when models are forced to select between US and Ghana correct answers, showing a bias towards US correct answers in unspecified contexts. While LLMs improve in specified cultural settings, encoder models still struggle with Ghanaian contexts, indicating a better attunement to US culture, likely due to the predominance of US-centric training data. The lower performance in Ghanaian specified settings, particularly among encoder models, underscores the need for better model training to handle non-US cultures.
5.3 Cultural Context Improves Answer Selection Accuracy
When models are presented with our test set questions that have specific cultural references for Ghana and US in the All-answers, Ghana-only and US-only configurations, we observe a trend where Llama3-70B always records the highest performance.
In the Ghana-only configuration, when forced to select between the correct and incorrect answers, Llama3-70B shows the highest accuracy of 90.67% in Figure 11, while in the US-only configuration, Llama3-70B achieves 95.62% in the US Specified setting, Figure 12. Roberta-base shows the highest performance for the encoder models with 69.71% for Ghana specified and 80% in the US Specified. These experiment results suggests that some of these models can effectively distinguish between country-specific correct answers and distractors, indicating some degree of cultural adaptability.
Figure  LABEL:fig:specified_accuracy shows the results in the All-answers configuration where we observe for the Ghana Specified setting, a lower preference for selecting the Ghana Correct Answer compared to the preference for selecting US Correct Answers in the US Specified setting. Ideally, we would expect the orange segments, fill up the full scale in the Ghana specified cultural setting, and the blue segments in the US specified cultural setting. While this ideal scenario is not fully realized, we still observe a higher accuracy in the US specified, with most models achieving above 50%. In the Ghana specified, Llama3-70B scored 60% and Gemma-7B scored 50% while all other models scored below 50%. Comparing this performance to the US specified, where we see Llama3-70B scoring 77% we realize a performance dsiscrepency. For most of these models, a combination of the correct and distractor answers exceeds the halfway mark which suggests that when some culture cues are provided, these models prefer selecting the specified cultural answer more than half the time.
We also observe this when moving from the unspecified to the specified settings, where we see a slight increase in the correct answer selection in Figure  LABEL:fig:specified_accuracy. Thus, embedding some cultural cues in the question, to an extent can guide the model towards selecting the correct cultural answer. This suggests that these models are able to some extent culturally adapt based on the cultural cues provided, indicated in their recorded accuracies. Notwithstanding this, there is still room for improvement in their cultural adaptability and preference, especially in the Ghanaian cultural settings.
5.4 Qualitative Analysis of Model Responses
Context: This person is married with two little kids. Question: How can Kojo make Bronya feel more magical for his family? Options: A. Decorate the Christmas tree with lots of presents underneath it, and decorate the house and the front porch with ornaments and lights B. Take them to the North Pole C. Ensure there is enough food, drinks and fun games D. Leave a couple of bite marks in the cookies that his children leave for Santa Correct Answer: Ensure there is enough food, drinks and fun games
We observe for the qualitative example above, the model responses in Appendix, D). We see that although some of the models can do a decent job, selecting the correct answers from the distractors, some of the models’ prefer the US correct answer more.
From these responses, both bert-base and RoBERTA-base selected the US correct answer, which aligns with a traditional Western Christmas celebration. Llama2-7B and Gemma-7B selected the answer, which aligns with the Ghanaian correct answer, that is, ensuring there is enough food, drinks, and fun games, and provide explanations which support their responses. Their response, indictae some level of familiarity with the Ghanaian culture and understanding that although While Llama2-13B, Llama3-70B and Mistral-7B also favored the US correct answers with explanations more consistent with US culture. The analysis underscores the dominance of Western cultural norms in model training, which can lead to biased and less relevant responses for non-Western contexts.
Conclusion: These findings emphasize the need for ongoing efforts to train models with culturally diverse datasets and involve human annotators from varied backgrounds. Future work should aim to reduce biases and enhance the models’ cultural competence across a broader range of cultural settings.
6 Related Work
Data Collection and Annotation
Researchers often rely on collaborative efforts from diverse cultures at different stages in their work. Huang et al. (2019) and Clark et al. (2020) allow diverse annotators to create commonsense questions, ensuring real-world relevance and promoting natural question generation. Ponti et al. (2020) strategically recruits translators to make cultural and linguistic adjustments, ensuring high-quality translations. Our work builds on this to allow broader participation, benefiting from diverse multi-cultural perspectives. While participants do not explicitly create the initial questions, they have the freedom to suggest refinements to questions and contexts, ensuring cultural accuracy and relevance. Zellers et al. (2018) uses adversarial methods to generate multiple-choice answers validated by crowd workers. Yin et al. (2021) employs human annotators for questions and correct answers, with automated distractors generation. Our approach involves annotators creating, selecting, and validating answers.
Cultural Sensitivity and Bias Detection in Commonsense Reasoning
Palta and Rudinger (2023), Zhou et al. (2023), Huang and Yang (2023), Bauer et al. (2023) and Naous et al. (2024) evaluate reasoning through downstream QA tasks, inference, entity recognition and sentiment analysis tasks, and discuss Western cultural biases. Shi et al. (2024) fine-tune LLMs on a dataset of self-narratives from TikTok, and measure cultural awareness through automatic entailment scores and human evaluations. Li et al. (2024) prompts models with culturally relevant sentences and analyses the text for cultural symbols and biases, focusing on diversity and implicit Western biases. Rao et al. (2024) assesses LLM responses to social acceptability questions across cultures, revealing difficulties with non-Western norms. Our study evaluates cultural adaptability with our test set that contains both explicit and implicit cultural references. We assess how well models can interpret and respond to culturally nuanced questions in a MCQ task.
7 Conclusion
We create a new test set to investigate cultural biases in models, focusing on Ghana and the US. The focus on the Ghanaian population, a harder-to-study yet important group, underlines the importance of representing diverse cultures in computational research. Our study offers multi-stage annotator participation throughout the dataset creation, ensuring we capture nuanced cultural specificities. Our experiments show that models favor US culture in unspecified cultural contexts and specifying cultural cues improves accuracy for both groups, albeit, the improvement is more pronounced for the US. These findings highlight the need for ongoing efforts to make models more culturally inclusive, particularly for low-resource languages. Mitigation techniques should involve annotators from diverse backgrounds at all stages of dataset creation, and training models tailored to specific cultures.
Limitations
Limited dimensions of culture Culture as a social concept, spans a wide range of practices, beliefs, and expressions. Our research focuses on specific cultural facets for only two select cultural groups, which represents just a small aspect of culture. This focus was necessary due to the scope of this study and the resource constraints our study presents. Future research could expand on this by exploring more cultural dimensions for more low-resource cultures, that will offer a more comprehensive understanding of diverse human cultures in language models.
Cultural stereotype There is a risk that some survey questions inadvertently perpetuate cultural stereotypes, particularly towards some ethnic groups, regions, or states. For instance, questions that reference a specific aspect of an ethnic group’s food culture, or social custom, that is well-known among other people in the country could create an implicit marker that primes participants to select a correct cultural context answer but potentially reinforce stereotypes about the ethnic group.
Incomplete representation of ethnic/region demographics The survey might overemphasize the opinions of the majority ethnic groups. This could occur if minority viewpoints are less recognized, leading to the majority’s perspective being perceived as the “correct” answer. As such, our best-voted answers might represent the viewpoints of the majority group while other good but least-voted answers although relatable to other groups are in the minority.
Limited dataset size and skewed demographics The size of our dataset is limited because we had to rely on volunteer annotators, as we could not use standard crowdsourcing tools for the Ghanaian population. This population is harder to study but important to understand due to its unique cultural context. Additionally, the participants we recruited skewed towards younger subjects, which may not fully represent the broader demographic distribution of the populations studied. This age skew could influence the cultural insights and relevance of the responses collected. Future research should aim to include a more diverse age range and employ strategies to increase participation from underrepresented groups.
Ethics Statement
Our research investigates cultural biases in commonsense reasoning models by exploring specific cultural facets. We recognize that exploring cultural nuances presents inherent risks, particularly in inadvertently simplifying or misrepresenting cultures. Our work which focused on two select cultures, inherently limits the breadth of cultural diversity and understanding. We involved annotators from diverse backgrounds from these two groups to validate our findings, but we acknowledge that this does not fully encompass the rich diversity within each culture. We also emphasize that while our findings may highlight the presence of cultural biases in certain models, it is crucial to approach the interpretation of our results with an understanding of their contextual limitations and the complexities of cultural representation in AI. Finally, we utilized AI assistants in our experiments (e.g., paraphrasing prompts, etc.).
Acknowledgements
We thank the anonymous reviewers for their constructive feedback. We also thank Kwesi Cobbina for their invaluable insights, as well as Nelson Padua-Perez, Julius Amegadzie, Belinda Laryea, Deborah Laryea, Amanda Acquaye and Gatha Adhikari for their support in recruiting annotators. We extend our heartfelt gratitude to the volunteers from Academic City University College, the University of Maryland, as well as all other volunteers from other institutions who contributed to this work. Their support was instrumental in the successful completion of our project. This work was supported by NSF CAREER Award No. 2339746. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Appendix A Dataset Statistics
Dataset No. of samples
CSQA 77
SIQA 53
PIQA 70
Table 3: Information showing the number of samples used from each dataset
Disambiguation Type No. of samples
Explicit 200
Implicit 150
Unspecified 175
Table 4: Information showing the number of samples for each disambiguation type. Unspecified types do not have any cultural reference(either overt or subtle).
Cultural Facet No. of samples
Social Customs and Lifestyle 354
Architecture/Housing 57
Food and Drinks 48
Gender, Marriage and Family 39
Geography 27
Total 525
Table 5: Test set breakdown for different cultural facets.
Appendix B Prompts
Three prompts were used, increasing in specificity and detail from Prompt 1 to Prompt 3, referencing specific contexts and settings.
B.1 Prompt 1
"You will be tasked with answering Multiple Choice Questions (MCQs). For each question, identify the answer that best fits the given context. Despite multiple options potentially being correct, select the one that is most appropriate for the specific scenario presented. Your response should be formatted in JSON, indicating the ’Best Answer’ along with its corresponding option number and text."
Figure 8: General MCQ task with no specific context provided.
B.2 Prompt 2
"You will be tasked with answering Multiple Choice Questions (MCQs). For each question, identify the answer that best fits the given context for a/an [specific setting] setting. Your response should be formatted in JSON, indicating the ’Best Answer’ along with its corresponding option number and text."
Figure 9: MCQ task with a specific setting provided.
B.3 Prompt 3
"You will be tasked with answering Multiple Choice Questions (MCQs). For each question, answer as a/an [specific culture] the answer that best fits the given context for a/an [specific setting]. Your response should be formatted in JSON, indicating the ’Best Answer’ along with its corresponding option number and text."
Figure 10: MCQ task with both a specific cultural perspective and setting provided.
Appendix C Additional Results
[…]
%GH %US %GH Dist. %US Dist. %GH %US %GH Dist. %US Dist. %GH %US %GH Dist. %US Dist.
bert-base 31.43 32.57 17.14 18.86 32.00 45.14 15.43 7.43 +0.57 +12.57 -1.71 -11.43
roberta-base 25.14 35.43 18.29 21.14 30.29 51.43 10.28 8.00 +5.15 +16.00 -8.01 -13.14
Llama2-7B 29.86 24.91 24.91 20.34 36.77 44.38 10.65 8.20 +6.91 +19.47 -14.26 -12.14
Llama2-13B 32.62 30.90 21.19 15.29 30.05 51.96 10.05 7.94 -2.57 +21.06 -11.14 -7.35
Llama2-70B 40.47 39.14 10.39 10.00 29.95 59.48 6.34 4.23 -10.52 +20.34 -4.05 -5.77
Llama3-70B 41.77 41.58 9.95 6.70 23.43 68.57 5.33 2.67 -18.34 +26.99 -4.62 -4.03
Gemma-7B 34.86 33.34 16.57 15.23 27.95 57.09 7.19 7.77 -6.91 +23.75 -9.38 -7.46
Mistral-7B 38.08 35.99 11.16 14.77 30.66 60.19 4.00 5.14 -7.42 +24.20 -7.16 -9.63
[…]
GH Specified US Specified
GH %US %GH Dist. %US Dist. GH %US %GH Dist. %US Dist.
bert-base 34.0 37.0 18.0 11.0 28.0 44.0 15.0 13.0
roberta-base 33.0 46.0 14.0 7.0 31.0 52.0 10.0 7.0
Llama2-7B 38.0 37.0 15.0 10.0 26.0 53.0 10.0 11.0
Llama2-13B 42.25 37.25 9.25 11.25 29.0 54.0 7.5 9.5
Llama2-70B 55.0 35.0 6.0 4.0 23.0 67.0 4.0 6.0
Llama3-70B 60.0 33.0 4.1 2.9 18.0 77.0 3.0 2.0
Gemma-7B 50.0 34.25 8.75 7.0 24.25 62.0 5.0 8.75
Mistral-7B 48.0 42.0 5.0 5.0 25.0 65.0 5.0 5.0
Table 7: Full Results on Model Preference Distribution in GH and US Specified Settings. Columns "%GH" and "%US" show the percentage of correct Ghanaian and US answers, "%GH Dist." the percentage of Ghanaian distractors, and "%US Dist." the percentage of US distractors.
UN Specified GH Specified US Specified
%GH %US %GH %US %GH %US
bert-base 41.0 59.0 47.0 53.0 41.0 59.0
roberta-base 37.0 63.0 42.0 58.0 39.0 61.0
Llama2-7B 51.0 49.0 49.0 51.0 37.0 63.0
Llama2-13B 35.0 65.0 58.0 42.0 31.0 69.0
Llama2-70B 37.0 63.0 66.0 34.0 25.0 75.0
Llama3-70B 26.0 74.0 67.0 33.0 18.0 82.0
Gemma-7B 40.0 60.0 59.0 41.0 30.0 70.0
Mistral-7B 34.0 66.0 56.0 44.0 25.0 75.0
Table 8: Full Results on Model Preference Distribution in UN, GH, and US Specified Settings for the Correct-only configuration. Columns "%GH" and "%US" show the percentage of correct Ghanaian and US answers.
Figure 11: Accuracy of models in Ghanaian settings when conditioned only on Ghanaian correct answers and Ghanaian distractors.
Figure 12: Accuracy of models in US settings when conditioned only on US correct answers and US distractors.
Appendix D Qualitative Results
This section shows the different model responses for one GH-specified test question when all four answer choices are provided.
Model Responses:
bert-base: Decorate the Christmas tree with lots of presents underneath it, and decorate the house and the front porch with ornaments and lights
roberta-base: Decorate the Christmas tree with lots of presents underneath it, and decorate the house and the front porch with ornaments and lights
Llama2-7B: "Best Answer": "(3) Ensure there is enough food, drinks, and fun games for the whole family.",
"Explanation": "The context of the question is that Kojo wants to make Bronya feel more magical for his family, and the options provided are all related to Christmas celebrations. Option (3) is the best answer because it suggests ensuring that there is enough food, drinks, and fun games for the whole family, which is a great way to create a magical and enjoyable experience for everyone."
Llama2-13B: "Best Answer": "(1) Decorate the Christmas tree with lots of presents underneath it, and decorate the house and the front porch with ornaments and lights",
"Explanation": "To make Bronya feel more magical for his family, Kojo should focus on creating a festive and joyful atmosphere around the Christmas tree. Decorating the tree with lots of presents underneath it, and the house and front porch with ornaments and lights, will help create a sense of wonder and excitement for the children. This option aligns best with the context of the question, which is focused on creating a magical atmosphere for the family during the holiday season."
Llama3-70B: "Best Answer": "(1) Decorate the Christmas tree with lots of presents underneath it, and decorate the house and the front porch with ornaments and lights",
"Explanation": "This answer is the most suitable because it creates a festive and magical atmosphere for Kojo’s family, especially his children, a key aspect of the Christmas season."
Gemma-7B: "Best Answer": "(3) Ensure there is enough food, drinks and fun games",
"Explanation": "The text describes a person who is married with two little kids and wants to make Bronya feel more magical for their family."
Mistral-7B: "Best Answer": "1. Decorate the Christmas tree with lots of presents underneath it, and decorate the house and the front porch with ornaments and lights.",
"Explanation": "The context mentions that Kojo has a family with two little kids. The question asks how Kojo can make Bronya feel more magical for his family. Option 1 is the best answer as it aligns with the context and the holiday season. Decorating the Christmas tree."
Appendix E Additional Experimental Setup Details
Encoder Models
We finetuned the encoder models using a consolidated version of the original datasets,(CSQA, SIQA, and PIQA) for 3 epochs and a learning rate of 1e-5 for the base models. The consolidated dataset for the models was created by standardizing the number of answer choices across the different datasets to four answer choices per question. We evaluated these finetuned encoder models on the validation sets of these original datasets to ensure it is consistent with existing baselines. Performance-wise, on the consolidated dataset, BERT-base achieved scores of 60.77% on CSQA, 60.29% on SIQA, and 65.83% on PIQA. Roberta-base performed better with scores of 69.37% on CSQA, 70.37% on SIQA, and 69.37% on PIQA.
LLMs
For the Llama2 models, Gemma and Mistral, we run the experiments with a random seed of 42. We used a temperature of 0.1 for all LLMs.
Summary of Experiments
Exp. Setting Cultural Setting Exp. Configuration Answers Provided
Unspecified No culture All-answers GH Correct, GH Dist., US Correct, US Dist.
No culture Correct-only GH Correct, US Correct
Specified GH All-answers GH Correct, GH Dist., US Correct, US Dist.
GH Correct-only GH Correct, US Correct
GH GH-only GH Correct, GH Dist.
Specified US All-answers GH Correct, GH Dist., US Correct, US Dist.
US Correct-only GH Correct, US Correct
US US-only US Correct, US Dist.
Answer-Only No culture All-answers GH Correct, GH Dist., US Correct, US Dist.
Table 9: This table summarizes all experiments we conduct. GH = Ghana, Dist. = Distractors.
Terms of use for each model
We carefully follow the guidelines per the terms of usage described by the model authors or company
Licenses
The CSQA, SIQA and PIQA are used under the MIT [^9] https://opensource.org/license/MIT and CC-BY [^10] https://creativecommons.org/licenses/by/4.0/ licenses.
Appendix F Disambiguation for the three datasets
Figure 13: An example of the question-answer format in the original CSQA Talmor et al. (2019) dataset and the disambiguated version from our dataset, with Ghanaian and American versions as well as the correct and distractor answers for both versions.
Figure 14: An example of the question-answer format in the original SIQA Sap et al. (2019) dataset and the disambiguated version from our dataset, with Ghanaian and American versions as well as the correct and distractor answers for both versions.
Figure 15: An example of the question-answer format in the original PIQA Bisk et al. (2020) dataset and the disambiguated version from our dataset, with Ghanaian and American versions as well as the correct and distractor answers for both versions.
Appendix G Surveys
Figure 16: Survey Sample from the Answer Choice Generation Stage for Ghana
Figure 17: Survey Sample from the Answer Choice Generation Stage for the USA
Figure 18: Survey Sample from the Likert Scale Answer Annotation Stage for Ghana
Figure 19: Survey Sample from the Likert Scale Answer Annotation Stage for the USA
Figure 20: Survey Sample from the Multiple Answer Choice Annotation for Ghana. The slider scales for the confidence ratings were locked to integers, ensuring that respondents provided answers as whole numbers only.
Figure 21: Survey Sample from the Multiple Answer Choice Annotation for the USA. The slider scales for the confidence ratings were locked to integers, ensuring that respondents provided answers as whole numbers only.
Appendix H Annotators
The annotators for each stage played a crucial role in validating the accuracy and relevance of these disambiguated questions and answers.
H.1 Annotator Elligibilty
We recruited participants who were adults aged 18 years or older, fluent in English and self-identified as either American or Ghanaian. The eligibility criteria further inquired about the nationality (Ghanaian/American), current country of residence (Ghana/USA), and place of most time spent before turning 18 (Ghana/USA). These criteria ensured that participants were well situated and could provide culturally relevant responses, with annotators from the United States focusing on a U.S. cultural context and those from Ghana on a Ghanaian cultural context.
H.2 Recruitment Process
Interested participants were invited to email their interest to a designated study-specific email address after which we present the full consent detailing the risks involved as well as other information relevant to the study. We responded to this interest email with a Qualtrics link for a particular survey set. We also obtained IRB approval for collecting, storing, and deleting the email addresses of participants. These email addresses were collected from direct email communications initiated by the participants themselves, who voluntarily reached out to express their interest in joining our study. The emails were used to track which survey set was assigned to each volunteer, ensuring that no participant received the same set in the same or different stages of the study. We ensure that all participants give their consent to this by presenting it as the first question in our survey. We anonymize all responses before processing so the emails are not tracked in our dataset creation process.
This rigorous recruitment process ensured that our participants were adequately informed and consented to participate, aligning with ethical research standards.
H.3 Annotator Demographics
Across all plots for both Ghanaian and American participants, we observe more participations in the Multiple Annotation stages. We present detailed demographic plots of our participants.These figures show some differences between Ghanaian and American annotators in terms of gender, age, education, ethnicity, and regional representation. Some observations include a higher number of male participants in the U.S. and a higher gender balance for Ghanaian annotators, with significant participation from younger individuals and varying educational backgrounds.
The ethnic group distribution for Ghanaians show more representation from the Akan and Ga-Dangme groups across all stages, while the distribution for Americans show more Asian and White participants across all stages. For detailed ethnicity breakdown, please refer to Figure 25.
Additionally, we see the regional distrbiton shows more representation from the Greater Accra Region across all stages, especially in the Multiple Answer Choice Annotation stage. For the American participants we also see more representation from the Northeast Region across all stages. Please refer to Figure 26.
Figure 22: Gender Distribution across Stages for Ghanaian and American Participants
Figure 23: Age Distribution across Stages for Ghanaian and American Participants
Figure 24: Education Distribution across Stages for Ghanaian and American Participants
Figure 25: Ethnicity Distribution across Stages for Ghanaian and American Participants
Figure 26: Region Distribution across Stages for Ghanaian and American Participants

### v2

@@
 These speculations highlight the importance of balanced training datasets to improve model performance across different cultural facets.
-5 Results and Discussion
-We analyse the performance of our selected models across our experimental setups to understand their inherent biases and cultural sensitivities.
-We present the models’ preferences in an unspecified cultural setting for the All-answers configuration in Table 1 and their overall preference in their respective specified cultural settings in Figure LABEL:fig:specified_accuracy.
-We also show findings for our specified experiments in our Correct-only configuration in Figure 7 as well as Ghana-only in Figures 11 and US-only in Figure 12 of the appendix.
-We also discuss qualitatively the model’s responses for a specific test set question and the implications of these realized trends.
-5.1 Inherent Preference For US Answers in Cultural Unspecified Contexts
-To establish a baseline, we run the unspecified experiment using the All-answers configuration and show their preference in the Answer-Only Section of Table 1.
-We observe that Roberta-base in this experiment shows a high propensity for selecting US correct answers while all other models do not record any significant differences in selecting US correct and Ghana correct answers.
-The relatively equal rates of selecting US correct and Ghana correct answers in the Answer-Only setup, show that there is nothing intrinsic about the Ghana correct answers that the model dislikes, e.g w.r.t phrasing, there is nothing superficially "wrong" with the Ghana correct answers.
-Compared to the Answer-Only setup, when a question with no cultural reference to either Ghana or the US is provided in the All-answers configuration, it reduces the chances of the models selecting distractor answers, indicated by the negative differences recorded in the ($\Delta$).
-This suggests that providing a question helps models pick up relevant cues to select correct answers, thereby reducing their preference for distractors.
-The positive differences for %US in this ($\Delta$) column and the higher rate of change indicate a higher preference for US correct answers compared to Ghana correct answers across all models, especially RoBERTa-base, which consistently prefers more US correct answers both in the Answer-Only as well as the Question-and-Answers.
-Llama3-70B shows the highest preference for US correct answers in this unspecified cultural setting.
-These results suggest that the models are more attuned to US culture, likely stemming from the predominance of US based data used in the pre-training of these models.
-5.2 Performance Discrepancy Favoring US
-For the unspecified cultural setting in the Correct-only configuration, the models prefer more US correct answers as shown in Figure 7 compared to Ghana correct answers which is always less than 50%.
-For the specified settings in the Correct-Only configuration, all LLMs prefer the specified cultural correct answer more than 50% which shows that the models have some ability to adapt to the specified cultural context.
-We observe for the Ghana specified, the encoder models prefer the Ghana correct answers less than 50% of the time.
-We also observe a higher preference of US correct answers in the US specified setting compared to the Ghana specified setting.
-Llama3-70B, the highest performing model in the Ghana specified in the Correct-only configuration, prefers Ghana correct answers 67% of the time while in the US Specified, it prefers US Correct Answers, 82% of the time.
-These findings highlight a performance discrepancy when models are forced to select between US and Ghana correct answers, showing a bias towards US correct answers in unspecified contexts.
-While LLMs improve in specified cultural settings, encoder models still struggle with Ghanaian contexts, indicating a better attunement to US culture, likely due to the predominance of US-centric training data.
-The lower performance in Ghanaian specified settings, particularly among encoder models, underscores the need for better model training to handle non-US cultures.
-5.3 Cultural Context Improves Answer Selection Accuracy
-When models are presented with our test set questions that have specific cultural references for Ghana and US in the All-answers, Ghana-only and US-only configurations, we observe a trend where Llama3-70B always records the highest performance.
-In the Ghana-only configuration, when forced to select between the correct and incorrect answers, Llama3-70B shows the highest accuracy of 90.67% in Figure 11, while in the US-only configuration, Llama3-70B achieves 95.62% in the US Specified setting, Figure 12.
-Roberta-base shows the highest performance for the encoder models with 69.71% for Ghana specified and 80% in the US Specified.
-These experiment results suggests that some of these models can effectively distinguish between country-specific correct answers and distractors, indicating some degree of cultural adaptability.
-Figure  LABEL:fig:specified_accuracy shows the results in the All-answers configuration where we observe for the Ghana Specified setting, a lower preference for selecting the Ghana Correct Answer compared to the preference for selecting US Correct Answers in the US Specified setting.
-Ideally, we would expect the orange segments, fill up the full scale in the Ghana specified cultural setting, and the blue segments in the US specified cultural setting.
-While this ideal scenario is not fully realized, we still observe a higher accuracy in the US specified, with most models achieving above 50%.
-In the Ghana specified, Llama3-70B scored 60% and Gemma-7B scored 50% while all other models scored below 50%.
-Comparing this performance to the US specified, where we see Llama3-70B scoring 77% we realize a performance dsiscrepency.
-For most of these models, a combination of the correct and distractor answers exceeds the halfway mark which suggests that when some culture cues are provided, these models prefer selecting the specified cultural answer more than half the time.
-We also observe this when moving from the unspecified to the specified settings, where we see a slight increase in the correct answer selection in Figure  LABEL:fig:specified_accuracy.
-Thus, embedding some cultural cues in the question, to an extent can guide the model towards selecting the correct cultural answer.
-This suggests that these models are able to some extent culturally adapt based on the cultural cues provided, indicated in their recorded accuracies.
-Notwithstanding this, there is still room for improvement in their cultural adaptability and preference, especially in the Ghanaian cultural settings.
-5.4 Qualitative Analysis of Model Responses
-Context: This person is married with two little kids.
-Question: How can Kojo make Bronya feel more magical for his family?
-Options: A.
-Decorate the Christmas tree with lots of presents underneath it, and decorate the house and the front porch with ornaments and lights B.
-Take them to the North Pole C.
-Ensure there is enough food, drinks and fun games D.
-Leave a couple of bite marks in the cookies that his children leave for Santa Correct Answer: Ensure there is enough food, drinks and fun games
-We observe for the qualitative example above, the model responses in Appendix, D).
-We see that although some of the models can do a decent job, selecting the correct answers from the distractors, some of the models’ prefer the US correct answer more.
-From these responses, both bert-base and RoBERTA-base selected the US correct answer, which aligns with a traditional Western Christmas celebration.
-Llama2-7B and Gemma-7B selected the answer, which aligns with the Ghanaian correct answer, that is, ensuring there is enough food, drinks, and fun games, and provide explanations which support their responses.
-Their response, indictae some level of familiarity with the Ghanaian culture and understanding that although While Llama2-13B, Llama3-70B and Mistral-7B also favored the US correct answers with explanations more consistent with US culture.
-The analysis underscores the dominance of Western cultural norms in model training, which can lead to biased and less relevant responses for non-Western contexts.
-Conclusion: These findings emphasize the need for ongoing efforts to train models with culturally diverse datasets and involve human annotators from varied backgrounds.
-Future work should aim to reduce biases and enhance the models’ cultural competence across a broader range of cultural settings.
 5 Related Work